OpenStudy (anonymous):

Factorize y=2x^2+3x-2

OpenStudy (anonymous):

+2 could be 1*2 or -1*-2 -2 could be -1*2 or 1*-2 Try all these combinations until you get the one that works, it's trial and error

OpenStudy (anonymous):

for instance (2x+1)(1x-2) foil it out and see what you get

OpenStudy (anonymous):

you get 2x^2+x-4x-2 = 2x^2-3x-2, not the same as the problem so that's not the right one

OpenStudy (anonymous):

(-2x+1)(-x-2) = 2x^2 -x +4x -2 = 2x^2 +3x-2, same as the problem, so (-2x+1)(-x-2) is the answer

OpenStudy (anonymous):

thankyouverymuch